OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16

hartnn (hartnn):

start by dividing both sides by -4 what u get ?

OpenStudy (anonymous):

. . .

hartnn (hartnn):

those are 3 dots if i am not wrong

hartnn (hartnn):

|x+5| = (-16)/(-4) = ... ?

OpenStudy (anonymous):

after you divide both sides by -4 you get: x+5=4 or x+5=-4 x= -1 or x= -9 and both solutions are correct because they give you 16 and -16.